# A primer on quasi-convex functions in nonlinear potential theories

### v2

@@ 7. Basic tools in nonlinear potential theory
    Let $X\subset{\mathbb{R}}^{n}$ be open. For any subequation $\mathcal{F}$ on $X$, the following properties hold:
    • (i:
      local)  $u\in\operatorname{USC}(X)$ locally $\mathcal{F}$-subharmonic $\iff$ $u\in\mathcal{F}(X)$;
    • (ii:
      maximum)  $u,v\in\mathcal{F}(X)$implies $\max\{u,v\}\in\mathcal{F}(X)$;
    • (iii:
      coherence)  if $u\in\operatorname{USC}(X)$ is twice differentiable at $x_{0}\in X$, then
@@
    • (iv:
      sliding)  $u\in\mathcal{F}(X)$implies $u-m\in\mathcal{F}(X)$ for any $m>0$;
    • (v:
      decreasing sequence)  $\{u_{k}\}_{k\in\mathbb{N}}\subset\mathcal{F}(X)$ decreasing $\implies$ $\lim_{k\to\infty}u_{k}\in\mathcal{F}(X)$;
    • (vi:
      uniform limits)  $\{u_{k}\}_{k\in\mathbb{N}}\subset\mathcal{F}(X)$, $u_{k}\to u$ locally uniformly $\implies$ $u\in\mathcal{F}(X)$;
    • (vii:
      families locally bounded above)  if $\mathscr{F}\subset\mathcal{F}(X)$ is a family of functions which are locally uniformly bounded above, then the upper semicontinuous envelope $u^{*}$ of the Perron function $u(\,\cdot\,)\vcentcolon=\sup_{w\in\mathscr{F}}w(\,\cdot\,)$ belongs to $\mathcal{F}(X)$. [^11] Recall that the upper semicontinuous envelope of a function $g$ is defined as the function $g^{\ast}(x)\vcentcolon=\lim_{r\searrow 0}\sup_{y\in B_{r}(x)}g(y).$ It is immediate to see that the upper semicontinuous envelope operator :*$*{}^{*}\colon g\mapsto g^{*}$ is the identity on the set of all upper semicontinuous functions. Also, we called Perron function the upper envelope of the family $\mathscr{F}$, since $\mathscr{F}$ is a family of subharmonics.
@@
    • (viii:
      translation)  $u\in\mathcal{F}(X)$ $\iff$ $u_{y}\vcentcolon=u(\cdot-y)\in\mathcal{F}(X+y)$, for any $y\in{\mathbb{R}}^{n}$.
@@
    (i). It follows immediately from the local nature of the notion of $\mathcal{F}$-subharmonicity.
    (ii). Let $w\vcentcolon=\max\{u,v\}$ and note that if $(p,A)\in J^{2,+}_{x}w$, then either $(p,A)\in J^{2,+}_{x}u$ or $(p,A)\in J^{2,+}_{x}v$ since $w(x)$ is either $u(x)$ or $v(x)$ and, in general, $u(y),v(y)\leqslant w(y)$. Therefore, in any case, $(w(x),p,A)\in\mathcal{F}_{x}$.
    (iii). It follows from the positivity condition (P), thanks to the elementary fact that, for $u$ which is at least twice differentiable at $x$, if $(p,A)$ is an upper contact jet for $u$ at $x$, then $(p,A)=(Du(x),D^{2}u(x)+P)$ for some $P\geqslant 0$.
    (iv). It follows immediately from the negativity condition (N).
    (v). We are going to use the definitional comparison. Since the decreasing limit $u$ of $\{u_{k}\}_{k\in\mathbb{N}}\subset\operatorname{USC}(X)$ belongs to $\operatorname{USC}(X)$, by Lemma 7.2 it suffices to show that for each $\Omega\Subset X$, one has $u+v\leqslant 0$ on $\Omega$ for each $v\in\operatorname{USC}(\overline{\Omega})\cap C^{2}(\Omega)$ which is strictly $\widetilde{\mathcal{F}}$-subharmonic on $\Omega$ with $u+v\leqslant 0$ on $\partial\Omega$. For each $\varepsilon>0$ fixed, consider the sets
@@
    By the upper semicontinuity of each $u_{k}$ (and thus of each $u_{k}+v$), $\mathfrak{E}_{k}$ are closed subsets of the compact set $\partial\Omega$ and hence each $\mathfrak{E}_{k}$ is compact. Since $\{u_{k}\}_{k\in\mathbb{N}}$ is decreasing, they are nested; that is, $\mathfrak{E}_{k+1}\subset\mathfrak{E}_{k}$ for all $k\in\mathbb{N}$. Since $u+v\leqslant 0$ on $\partial\Omega$, they have empty intersection:
@@
    Therefore, by the well-known property of nested families of compact sets, the $\mathfrak{E}_{k}$’s must be eventually empty. This means that $u_{k}+v<\varepsilon$ on $\partial\Omega$ for $k$ large and since $v-\varepsilon$ is still a smooth and strict $\widetilde{\mathcal{F}}$-subharmonic by the sliding property, the definitional comparison lemma yields $u_{k}+v\leqslant\varepsilon$ on $\Omega$. Hence $u+v\leqslant u_{k}+v\leqslant\varepsilon$ on $\Omega$, for any $\varepsilon>0$, which, on letting $\varepsilon\searrow 0$, gives $u+v\leqslant 0$ on $\Omega$ and thus $u\in\mathcal{F}(X)$, again by definitional comparison.
@@
    (vi). If $u_{k}$ converges to $u$ uniformly on the compact $\overline{\Omega}$ and we suppose $u+v\leqslant 0$ on $\partial\Omega$ as above, then, for any $\varepsilon>0$ fixed, $u_{k}+v<\varepsilon$ on $\partial\Omega$ for $k$ large. The definitional comparison lemma yields $u_{k}+v\leqslant\varepsilon$ on $\Omega$, hence $u+v=(u-u_{k})+(u_{k}+v)<2\varepsilon$ on $\Omega$ for $k$ large, where one exploits the uniform convergence to say that $u-u_{k}<\varepsilon$ on $\Omega$ if $k$ is large. On letting $\varepsilon\searrow 0$, we have $u+v\leqslant 0$ on $\Omega$ and thus, by the converse implication of the definitional comparison, $u\in\mathcal{F}(X)$.
@@
    (vii). If $u^{\ast}+v\leqslant 0$ on $\partial\Omega$, then $w+v\leqslant 0$ on $\partial\Omega$ for all $w\in\mathscr{F}$. By the implication (i) of the definitional comparison lemma, $w+v\leqslant 0$ on $\Omega$ for all $w\in\mathscr{F}$ and thus $u+v\leqslant 0$ on $\Omega$. Since $v$ is smooth and since the upper semicontinuous envelope operator preserves inequalities, we conclude that $u^{*}+v=(u+v)^{*}\leqslant 0$ on $\Omega$ and the thesis follows from the implication (ii) of definitional comparison.
@@
    (viii). This property holds only in a constant coefficient context; indeed, its validity for every $y\in{\mathbb{R}}^{n}$ is actually equivalent to the fiber $\mathcal{F}_{x}$ being independent of the point $x\in X$. Indeed, since by the respective definitions $u_{y}(x+y)=u(x)$ and $J^{2,+}_{x}u=J^{2,+}_{x+y}u_{y}$, one has, for $z\in X+y$, that $J^{2,+}_{z}u_{y}\subset\mathcal{F}_{z-y}$ and thus property (viii) holds if and only if $\mathcal{F}_{z-y}=\mathcal{F}_{z}$. ∎
    As a final remark, the coherence property (iii) says that an upper semicontinuous function which is twice differentiable and classically $\mathcal{F}$-subharmonic on $X$ (it satisfies the condition (5.14 for all $x\in X$), then it is $\mathcal{F}$-subharmonic in the viscosity sense. One might hope to have the same conclusion even if (5.14) is satisfied almost everywhere, with respect to the Lebesgue measure. This is true if $u$ is locally quasi-convex and is known as the Almost Everywhere Theorem from [26]. This deep result will be discussed in the next section. It was first established in the pure second-order case in [23, Corollary 7.5].
    8. Quasi-convex functions and subharmonics
    In this section, we treat the main theme of this work on the efficient use of quasi-convex function in nonlinear potential theories. We begin with a first fundamental result, which says that in the potential theory defined by a given subequation constraint set $\mathcal{F}$, a locally quasi-convex function is $\mathcal{F}$-subharmonic on an open set if it is $\mathcal{F}$-subharmonic on a set of full (Lebesgue) measure. Moreover, by Alexandrov’s Theorem 2.2 and the coherence property of Remark 5.3, the $\mathcal{F}$-subharmonicity then reduces to a classical statement almost everywhere.
@@ 8. Quasi-convex functions and subharmonics
    Suppose that $\mathcal{F}$ is a primitive subequation on an open set $X\subset{\mathbb{R}}^{n}$ and $u\colon X\to{\mathbb{R}}$ is a locally quasi-convex function. Then
@@
    By Remark 5.13 with $E\vcentcolon=\{x\in X:\ J^{2}_{x}u\in\mathcal{F}_{x}\}$, we know that $\mathcal{J}^{2,+}u\subset\overline{\mathcal{J}(u,E)}$, where $\mathcal{J}^{2,+}u$ is defined in (5.21) and $\mathcal{J}(u,E)$ is defined in (5.22); that is,
    $\mathcal{J}^{2,+}u:=\{(x,u(x),p,A):\ x\in X,(p,A)\in J^{2,+}_{x}u\}$
    and
    $\mathcal{J}(u,E)\vcentcolon=\big{\{}(x,u(x),Du(x),D^{2}u(x)+P):\ x\in E\cap\operatorname{Diff}^{2}(u),\ P\geqslant 0\big{\}}.$
    By property (P) one has $\mathcal{J}(u,E)\subset\mathcal{F}$ and by property (C) also $\overline{\mathcal{J}(u,E)}\subset\mathcal{F}$. Hence by Remark 5.12, $u\in\mathcal{F}(X)$.
    Conversely, for every $P>0$, we know that $(Du(x),D^{2}u(x)+P)$ is an upper contact jet for $u$ at $x\in\operatorname{Diff}^{2}(u)$. Therefore, in particular, $J^{2}_{x}u+(0,0,\varepsilon I)\in\mathcal{F}_{x}$ for all $\varepsilon>0$. By taking $\varepsilon\searrow 0$ and using the fact that $\mathcal{F}$ is closed, we get $J_{x}^{2}u\in\mathcal{F}_{x}$ for all $x\in\operatorname{Diff}^{2}(u)$ and hence the conclusion follows from Alexandrov’s theorem. ∎
@@
    The proof avove relies heavily on Remarks 5.12 and 5.13, which, in some sense, “hide the hard analysis” of the combined use of the Jensen–Słodkowski Theorem 3.28 and of Alexandrov’s Theorem 2.2. For somewhat more explicit proofs one can consult [26] or [8, Lemma 2.10], where the Bad Test Jet Lemma is also used.
    We are now ready for the second fundamental result. As noted by Harvey and Lawson in [26], an immediate consequence of the Almost Everywhere Theorem 8.1 is a subharmonic addition theorem for locally quasi-convex functions. That is, by restricting attention to suharmonics which are locally quasi-convex, then a purely algebraic statement of jet addition always implies the funcional analytic statement of subharmonic addition. This can be written as
    $\mathcal{F}_{x}+\mathcal{G}_{x}\subset\mathcal{H}_{x},\ \ \forall\,x\in X\implies(\mathcal{F}(X)\cap\mathsf{qc}_{\text{loc}}(X))+(\mathcal{G}(X)\cap\mathsf{qc}_{\text{loc}}(X))\subset\mathcal{H}(X),$ (8.1)
    where $\mathsf{qc}_{\text{loc}}(X)$ is the space of all locally quasi-convex functions on $X$, which were characterized in formula (5.26) of Example 5.18. The result is formalized in the following way.
    Theorem 8.3 (Subharmonic addition for quasi-convex functions).
    Suppose $\mathcal{F}$ and $\mathcal{G}$ are primitive subequations on an open set $X\subset{\mathbb{R}}^{n}$ and let $\mathcal{K}\vcentcolon=\overline{\mathcal{F}+\mathcal{G}}$; suppose $u\in\mathcal{F}(X)$ and $v\in\mathcal{G}(X)$ are locally quasi-convex functions on $X$. Then $u+v\in\mathcal{K}(X)$.
@@
    Since $u\in\mathcal{F}(X)$ and $v\in\mathcal{G}(X)$ are locally quasi-convex, by the reverse implication in the Almost Everywhere Theorem 8.1 we have $J^{2}_{x}u\in\mathcal{F}_{x}$ and $J^{2}_{x}v\in\mathcal{G}_{x}$ for almost every $x\in X$. Hence $J^{2}_{x}(u+v)=J^{2}_{x}u+J^{2}_{x}v\in\mathcal{K}_{x}$ almost everywhere, with $u+v$ locally quasi-convex. The forward implication of the Almost Everywhere Theorem 8.1 now yields $u+v\in\mathcal{K}(X)$. ∎
@@
    for some $\alpha\in(\frac{1}{3},\frac{2}{3})$. Then it is easy to see that $(0,0,0,0)\notin\mathcal{F}+\mathcal{G}$, yet it is a limit point, since for every positive integer $k$, we have that $\sqrt{1/3+k}-\sqrt{2/3-\alpha+k}=O\big{(}1/\sqrt{k}\big{)}$ belongs to the ${\mathbb{R}}$-component of each jet in $\mathcal{F}+\mathcal{G}$.
    With respect to Theorem 8.3 on subharmonic addition, two observations are in order. The first observation contains a reformulation of the Summand Theorem 2.16 which provides a stronger version of Theorem 8.3 while the second observation notes, as in [26], that the quasi-convexity assumption on the subharmonics is not needed in the constant coefficient case.
@@
    It is worth noting that the Summand Theorem 2.16 leads to a slightly stronger form of Theorem 8.3. Indeed, it tells us that if $(p,A)$ is an upper contact jet for the sum $w=u+v$ at $x$, then there exist two jets
@@
    which sum to $(w(x),p,A)$. To see this, note that the sequence $\{x_{j}\}$ can be selected in such a way that $Du(x_{j})\to Du(x)$ and $Dv(x_{j})\to Dv(x)$, according to Lemma 2.15; hence it is sufficient to take $B^{\prime}=A+P$, where $P=A-B-C\geqslant 0$. To write this succinctly,
@@
    where the former inclusion is trivial and the latter is the thesis of the proposition.
    In other words, Theorem 2.16 assures that every upper contact $2$-jet of the sum $u+v$ of two quasi-convex functions can be represented as the sum of two $2$-jets of $\mathcal{F}$ and $\mathcal{G}$, respectively, hence $u+v\in\mathcal{F}+\mathcal{G}\,(X)$. This will be used in the proof of the strict comparison for quasi-convex functions (Theorem 9.17).
@@
    Furthermore, if $\mathcal{F}$ and $\mathcal{G}$ are constant coefficients subequations, then the assumption that $u$ and $v$ are quasi-convex can be dropped; therefore we have
@@
    This follows from a quasi-convex approximation technique that was introduced in [23] in a context of pure second-order subequations and that works fine also for general constant coefficient subequations.
    As we anticipated in the previous part, the supconvolution is known to provide an effective way to approximate upper semicontinuous functions by a quasi-convex regularization of them (see Theorem 4.9 for four basic properties of the supconvolution). The next proposition which imitates [23, Theorem 8.2] shows that, for constant coefficent subequations $\mathcal{F}$, if one takes the supconvolution of a bounded $\mathcal{F}$-subharmonic with, then it remains $\mathcal{F}$-subharmonic, provided we are far enough from the boundary of the domain $X$.
    Proposition 8.8 (Quasi-convex approximation of $\mathcal{F}$-subharmonics).
    Let $\mathcal{F}$ be a constant coefficient subequation and suppose that $u\in\mathcal{F}(X)$ is bounded. For $\varepsilon>0$, let $\delta\vcentcolon=2\sqrt{\varepsilon\sup_{X}\!|u|}$, and $X_{\delta}\vcentcolon=\{y\in X:\ d(y,\partial X)>\delta\}$. Then $u^{\varepsilon}\in\mathcal{F}(X_{\delta})$.
@@
    Since $\mathcal{F}$ has constant coefficients, we have that $u_{z}\vcentcolon=u(\cdot-z)\in\mathcal{F}(X_{\delta})$ for all $z\in B_{\delta}$. Indeed, it is easy to see that $J^{2,+}_{x}u_{z}\subset\mathcal{F}_{x-z}=\mathcal{F}_{x}$ for every $x\in X_{\delta}$. If we now take
@@
    then $\mathscr{F}\subset\mathcal{F}(X_{\delta})$ by the negativity condition (N) and, by (4.8), its upper envelope is $u^{\varepsilon}$, and it coincides with $(u^{\varepsilon})^{\ast}$ because it is quasi-convex, hence upper semicontinuous. Therefore, we have $u^{\varepsilon}\in\mathcal{F}(X_{\delta})$, by the families-locally-bounded-above property of Proposition 7.4. ∎
    The final result of this section with an important application of the quasi-convex approximation technique by deducing a Subharmonic Addition Theorem for merely upper semicontinuous subharmonics for constant coefficient subequations from its quasi-convex counterpart Theorem 8.3. This result was first stated in [8, Theorem 7.1] and we essentially reproduce the original proof.
@@
    Suppose $\mathcal{F}$ and $\mathcal{G}$ are constant coefficient subequations. Then, for any open set $X\subset{\mathbb{R}}^{n}$, jet addition implies subharmonic addition; that is,

### v3

@@ 7.2. Fundamental properties of subharmonics
Let $X\subset{\mathbb{R}}^{n}$ be open, and let $\mathcal{F}$ be a subequation on $X$. Then the following properties hold:
  1. (i)
    local: given $u\in\operatorname{USC}(X)$,
@@
  2. (ii)
    maximum: $u,v\in\mathcal{F}(X)$implies $\max\{u,v\}\in\mathcal{F}(X)$;
  3. (iii)
    coherence: if $u\in\operatorname{USC}(X)$ is twice differentiable at $x_{0}\in X$, then
[…]
  4. (iv)
    sliding: $u\in\mathcal{F}(X)$implies $u-m\in\mathcal{F}(X)$ for any $m>0$;
  5. (v)
    decreasing sequences: given a (pointwise) decreasing sequence $\{u_{k}\}_{k\in\mathbb{N}}\subset\mathcal{F}(X)$, one has ${\lim_{k\to\infty}u_{k}\in\mathcal{F}(X)}$;
  6. (vi)
    uniform limits: if $\{u_{k}\}_{k\in\mathbb{N}}\subset\mathcal{F}(X)$ converges to $u$ locally uniformly on $X$, then $u\in\mathcal{F}(X)$;
  7. (vii)
    families locally bounded above: if $\mathscr{F}\subset\mathcal{F}(X)$ is a family of functions which are locally uniformly bounded above, and $u\vcentcolon=\sup_{w\in\mathscr{F}}w$ is the associated Perron function, then $u^{*}\in\mathcal{F}(X)$, where $u^{*}$ denotes the upper semicontinuous envelope of $u$ [^2] Recall that the upper semicontinuous envelope of a function $g$ is defined as the function $g^{\ast}(x)\vcentcolon=\lim_{r\searrow 0}\sup_{y\in B_{r}(x)}g(y).$ It is immediate to see that the upper semicontinuous envelope operator ${}^{*}\colon g\mapsto g^{*}$ is the identity on the set of all upper semicontinuous functions. Also, we called Perron function the upper envelope of the family $\mathscr{F}$, since $\mathscr{F}$ is a family of subharmonics.
@@
  1. (viii)
    translation: $u\in\mathcal{F}(X)$ $\iff$ $u_{y}\vcentcolon=u(\,\cdot-y)\in\mathcal{F}(X+y)$, for any $y\in{\mathbb{R}}^{n}$.
@@
[…]
To prove property (v), we are going to use the definitional comparison. Since the decreasing limit $u$ of $\{u_{k}\}_{k\in\mathbb{N}}\subset\operatorname{USC}(X)$ belongs to $\operatorname{USC}(X)$, by Lemma 7.3 it suffices to show that for each $\Omega\Subset X$, one has $u+v\leqslant 0$ on $\Omega$ for each $v\in\operatorname{USC}(\overline{\Omega})\cap C^{2}(\Omega)$ which is strictly $\widetilde{\mathcal{F}}$-subharmonic on $\Omega$ with $u+v\leqslant 0$ on $\partial\Omega$. For each $\varepsilon>0$ fixed, consider the sets
@@
By the upper semicontinuity of each $u_{k}$ (and thus of each $u_{k}+v$), $\mathfrak{E}_{k}$ are closed subsets of the compact set $\partial\Omega$ and hence each $\mathfrak{E}_{k}$ is compact; since $\{u_{k}\}_{k\in\mathbb{N}}$ is decreasing, they are nested (that is, $\mathfrak{E}_{k+1}\subset\mathfrak{E}_{k}$ for all $k\in\mathbb{N}$), and since $u+v\leqslant 0$ on $\partial\Omega$, they have empty intersection, which means that
@@
Therefore, by the well-known property of nested families of compact sets, $\mathfrak{E}_{k}$ must be empty for all large $k$. This means that $u_{k}+v<\varepsilon$ on $\partial\Omega$ for $k$ large and since $v-\varepsilon$ is still smooth and a strict $\widetilde{\mathcal{F}}$-subharmonic (by the sliding property (iv)), the definitional comparison lemma yields $u_{k}+v\leqslant\varepsilon$ on $\Omega$. Hence $u+v\leqslant u_{k}+v\leqslant\varepsilon$ on $\Omega$, for every $\varepsilon>0$. Taking the limit $\varepsilon\searrow 0$ gives $u+v\leqslant 0$ on $\Omega$ and thus $u\in\mathcal{F}(X)$, again by definitional comparison.
@@
As for property (vi), note that if $u_{k}$ converges to $u$ uniformly on the compact set $\overline{\Omega}$ and if we suppose $u+v\leqslant 0$ on $\partial\Omega$ as above, then, for any $\varepsilon>0$ fixed, $u_{k}+v<\varepsilon$ on $\partial\Omega$ for $k$ large. The definitional comparison lemma yields $u_{k}+v\leqslant\varepsilon$ on $\Omega$, hence $u+v=(u-u_{k})+(u_{k}+v)<2\varepsilon$ on $\Omega$ for $k$ large, where one exploits the uniform convergence to say that $u-u_{k}<\varepsilon$ on $\Omega$ if $k$ is large. Taking the limit $\varepsilon\searrow 0$ gives $u+v\leqslant 0$ on $\Omega$ and thus, by the converse implication of the definitional comparison, $u\in\mathcal{F}(X)$.
@@
Property (vii) is also proved by using the definitional comparison. If $u^{\ast}+v\leqslant 0$ on $\partial\Omega$, then $w+v\leqslant 0$ on $\partial\Omega$ for all $w\in\mathscr{F}$; then, by the implication (i) of the Definitional Comparison Lemma 7.3, $w+v\leqslant 0$ on $\Omega$ for all $w\in\mathscr{F}$ and thus $u+v\leqslant 0$ on $\Omega$. Since $v$ is smooth and since the upper semicontinuous envelope operator preserves inequalities, we conclude that $u^{*}+v=(u+v)^{*}\leqslant 0$ on $\Omega$ and the thesis follows from the implication (ii) of Lemma 7.3.
@@
Finally, we observe that property (viii) holds only in a constant coefficient context, as its validity for every $y\in{\mathbb{R}}^{n}$ is actually equivalent to the fiber $\mathcal{F}_{x}$ being independent of the point $x\in X$. Indeed, since by the respective definitions $u_{y}(x+y)=u(x)$ and $J^{2,+}_{x}u=J^{2,+}_{x+y}u_{y}$, one has, for $z\in X+y$, that $J^{2,+}_{z}u_{y}\subset\mathcal{F}_{z-y}$, and thus property (viii) holds if and only if $\mathcal{F}_{z-y}=\mathcal{F}_{z}$. ∎
[…]
Chapter 8 Semiconvex functions and subharmonics
[…]
8.1. The Almost Everywhere Theorem
The first essential result is the following theorem.
@@ 8.1. The Almost Everywhere Theorem
Suppose that $\mathcal{F}$ is a primitive subequation on an open set $X\subset{\mathbb{R}}^{n}$ and $u\colon X\to{\mathbb{R}}$ is a locally semiconvex function. Then
@@
By Remark 5.13 with $E\vcentcolon=\{x\in X:\,J^{2}_{x}u\in\mathcal{F}_{x}\}$, we know that
$\mathcal{J}^{2,+}u\subset\overline{\mathcal{J}(u,E)},$
where $\mathcal{J}^{2,+}u$ and $\mathcal{J}(u,E)$ are defined in (5.14) and (5.15), respectively. By property (P) one has $\mathcal{J}(u,E)\subset\mathcal{F}$ and by property (C) also $\overline{\mathcal{J}(u,E)}\subset\mathcal{F}$. Hence $u\in\mathcal{F}(X)$ by Remark 5.12.
Conversely, for every $P>0$, we know that $(Du(x),D^{2}u(x)+P)$ is an upper contact jet for $u$ at $x\in\operatorname{Diff}^{2}(u)$. Therefore, in particular, $J^{2}_{x}u+(0,0,\varepsilon I)\in\mathcal{F}_{x}$ for all $\varepsilon>0$. By taking $\varepsilon\searrow 0$ and using the fact that $\mathcal{F}$ is closed, we get $J_{x}^{2}u\in\mathcal{F}_{x}$ for all $x\in\operatorname{Diff}^{2}u$ and hence the conclusion follows from Alexandrov’s Theorem 2.2. ∎
@@
The proof above relies heavily on Remarks 5.12 and 5.13, which, in some sense, “hide the hard analysis” of the combined use of the Jensen–Słodkowski Theorem 3.27 and of Alexandrov’s Theorem 2.2. For somewhat more explicit proofs one can consult [29], or [9, Lemma 2.10] where the Bad Test Jet Lemma is also used.
8.2. The Subharmonic Addition Theorem
We are now ready for the second fundamental result. As noted in [29], an immediate consequence of the Almost Everywhere Theorem 8.1 is a subharmonic addition theorem for locally semiconvex functions. That is, by restricting attention to subharmonics which are locally semiconvex, a purely algebraic statement of jet addition always implies the functional analytic statement of subharmonic addition. This can be written as
$\mathcal{F}_{x}+\mathcal{G}_{x}\subset\mathcal{H}_{x}\ \ \forall\,x\in X\ \ \implies\ \ (\mathcal{F}(X)\cap\mathsf{sc}_{\mathrm{loc}}(X))+(\mathcal{G}(X)\cap\mathsf{sc}_{\mathrm{loc}}(X))\subset\mathcal{H}(X),$ (8.1)
where $\mathsf{sc}_{\mathrm{loc}}(X)$ is the space of all locally semiconvex functions on $X$, which were characterized in formula (5.16) of Example 5.18. The result is formalized in the following way.
Theorem 8.3 (Subharmonic addition for semiconvex functions).
Let $\mathcal{F}$ and $\mathcal{G}$ be primitive subequations on an open set $X\subset{\mathbb{R}}^{n}$ and define $\mathcal{K}\vcentcolon=\overline{\mathcal{F}+\mathcal{G}}$; let $u\in\mathcal{F}(X)$ and $v\in\mathcal{G}(X)$ be locally semiconvex functions on $X$. Then $u+v\in\mathcal{K}(X)$.
@@ 8.2. The Subharmonic Addition Theorem
Since $u\in\mathcal{F}(X)$ and $v\in\mathcal{G}(X)$ are locally semiconvex, by the reverse implication in the Almost Everywhere Theorem 8.1 we have $J^{2}_{x}u\in\mathcal{F}_{x}$ and $J^{2}_{x}v\in\mathcal{G}_{x}$ for almost every $x\in X$. Hence $J^{2}_{x}(u+v)=J^{2}_{x}u+J^{2}_{x}v\in\mathcal{K}_{x}$ almost everywhere, with $u+v$ locally semiconvex. The forward implication of the Almost Everywhere Theorem 8.1 now yields $u+v\in\mathcal{K}(X)$. ∎
@@
for some $\alpha\in(\frac{1}{3},\frac{2}{3})$, where $\chi_{E}$ denotes the characteristic function of the set $E$ (that is, $\chi(r)=1$ if $r\in E$ and $\chi(r)=0$ if $r\notin E$). It is easy to see that $(0,0,0,0)\notin\mathcal{F}+\mathcal{G}$, yet it is a limit point, since for every positive integer $k$, we have that $\sqrt{1/3+k}-\sqrt{2/3-\alpha+k}=O\big{(}1/\sqrt{k}\big{)}$ belongs to the ${\mathbb{R}}$ component of each jet in $\mathcal{F}+\mathcal{G}$.
With respect to Theorem 8.3 on subharmonic addition, two observations are in order. The first observation contains a reformulation of the Summand Theorem 2.16 which provides a stronger version of Theorem 8.3 while the second observation notes, as in [29], that the semiconvexity assumption on the subharmonics is not needed in the constant coefficient case.
@@
The Summand Theorem 2.16 leads to a slightly stronger form of Theorem 8.3. Indeed, it tells us that if $(p,A)$ is an upper contact jet for the sum $w=u+v$ at $x$, then there exist two jets
@@
which sum to $(w(x),p,A)$. To see this, note that, considering the sequence $\{x_{j}\}_{j\in\mathbb{N}}$ in Theorem 2.16(ii), it is sufficient to take $B^{\prime}=A+P$, where $P=A-B-C\geqslant 0$, or, more succinctly, note that
@@
where the former inclusion is trivial and the latter is given by Theorem 2.16.
In other words, Theorem 2.16 assures that every upper contact $2$-jet of the sum $u+v$ of two semiconvex functions $u\in\mathcal{F}(X)$ and $v\in\mathcal{G}(X)$ can be represented as the sum of two $2$-jets of $\mathcal{F}$ and $\mathcal{G}$, respectively, hence $u+v\in(\mathcal{F}+\mathcal{G})(X)$. This will be used in the proof of the strict comparison for semiconvex functions (Theorem 9.19).
@@
If $\mathcal{F}$ and $\mathcal{G}$ are constant coefficients subequations, then the assumption that $u$ and $v$ are semiconvex can be dropped; therefore, we have
@@
This will be proved next and follows from the semiconvex approximation technique that was introduced in [26] in a context of pure second-order subequations and that works fine also for general constant coefficient subequations.
[…]
@@
Since $\mathcal{F}$ is constant coefficient, we have that $u_{z}\vcentcolon=u(\cdot-z)\in\mathcal{F}(X_{\delta})$ for all $z\in B_{\delta}$. Indeed, it is easy to see that $J^{2,+}_{x}u_{z}\subset\mathcal{F}_{x-z}=\mathcal{F}_{x}$ for every $x\in X_{\delta}$. If we now take
@@
then $\mathscr{F}\subset\mathcal{F}(X_{\delta})$ by the negativity condition (N) and, by (4.6), its upper envelope is $u^{\varepsilon}$, and it coincides with $(u^{\varepsilon})^{\ast}$ because it is semiconvex, hence upper semicontinuous. Therefore, we have $u^{\varepsilon}\in\mathcal{F}(X_{\delta})$, by property (vii) of Proposition 7.5. ∎
The final result of this section is an important application of the semiconvex approximation technique. We are going to deduce a Subharmonic Addition Theorem for merely upper semicontinuous subharmonics for constant coefficient subequations from its semiconvex counterpart (Theorem 8.3). This was first stated in [9, Theorem 7.1] and we essentially reproduce the original proof.
@@
Let $\mathcal{F}$, $\mathcal{G}$ and $\mathcal{H}$ be constant coefficient subequations. Then, for any open set $X\subset{\mathbb{R}}^{n}$, jet addition implies subharmonic addition; that is,